# LLM Prompt Duel Optimizer: Efficient Label-Free Prompt Optimization

### v1

@@ Abstract
Large language models (LLMs) are highly sensitive to their input prompts, making prompt design a central challenge. While automatic prompt optimization (APO) reduces manual engineering, most approaches assume access to ground-truth references such as labeled validation data. In practice, however, collecting high-quality labels is costly and slow. We propose the Prompt Duel Optimizer (PDO), a sample-efficient framework for label-free prompt optimization. PDO formulates the problem as a dueling-bandit setting, where supervision signal comes from pairwise preference feedback provided by an LLM judge. The framework combines Double Thompson Sampling (D-TS), which prioritizes informative prompt comparisons, with Top-Performer Guided Mutation, which expands the candidate pool by mutating high-performing prompts. PDO naturally operates in label-free settings and can also incorporate partial labels to mitigate judge noise. Experiments on BIG-bench Hard (BBH) and MS MARCO show that PDO consistently outperforms baseline methods. Ablation studies further demonstrate the effectiveness of both D-TS and prompt mutation.
@@
Yuanchen Wu1thanks: Work done during an internship at Meta.thanks: Code will be available at https://github.com/meta-llama/prompt-ops   Saurabh Verma2   Justin Lee2   Fangzhou Xiong2 Poppy Zhang2Amel Awadelkarim2Xu Chen2Yubai Yuan1Shawndra Hill2 1Department of Statistics, The Pennsylvania State University 2Meta
@@ 1 Introduction
To address these issues, we propose Prompt Duel Optimizer (PDO), which treats preference-based prompt optimization as a dueling bandit problem. PDO uses Double Thompson Sampling (D-TS) (Wu and Liu, 2016) to adaptively choose which prompt pairs to compare. D-TS maintains posteriors over pairwise win probabilities and draws Thompson samples for both candidates, prioritizing informative comparisons. PDO also incorporates Top-Performer Guided Mutation, which periodically mutates top-performing prompts to generate local variants and prunes weaker ones. Aligning this expand-and-prune cycle with comparisons chosen by D-TS concentrates search on stronger regions of the prompt space without exhaustive enumeration.
@@
    We formulate preference-based prompt optimization without ground-truth label references as a dueling bandit problem with LLM-judged pairwise comparisons.
@@
    We introduce Prompt Duel Optimizer (PDO) and evaluate it on both multiple-choice and open-ended tasks. PDO outperforms heuristic and label-free baselines, and is also competitive with supervised APO methods.
@@
    Through ablation studies, we show: (i) D-TS is sample-efficient, identifying high-quality prompts with far fewer comparisons than random sampling; (ii) mutation steers search toward stronger regions and yields better candidates; and (iii) pairwise preference signals are more reliable than pointwise LLM scoring.
@@
    PDO flexibly incorporates a small fraction of ground-truth labels when available, mitigating judge noise, accelerating convergence, and enabling practical human-in-the-loop deployment.
@@ 3.1 Efficient Prompt Selection via Double Thompson Sampling
This two-step prompt selection procedure using D-TS achieves the following regret bounds in general Copeland settings.
Theorem 1 (PDO Copeland Regret)
Let $\mathcal{P}=\{p_{1},\dots,p_{K}\}$ be $K$ prompts with pairwise preference probabilities $\mu_{i,j}=\Pr(p_{i}\succ p_{j})$. Define the normalized Copeland score
$\zeta_{i}=\tfrac{1}{K-1}\sum_{j\neq i}\mathbf{1}\{\mu_{i,j}>1/2\},\quad\zeta^{*}=\max_{i}\zeta_{i}.$
If at round $t$ the algorithm selects $(a_{t},b_{t})$, the regret is $r_{t}=\zeta^{*}-\max\{\zeta_{a_{t}},\zeta_{b_{t}}\}$ and $R_{T}=\sum_{t=1}^{T}r_{t}$. Then Double Thompson Sampling (D-TS) achieves (Wu and Liu, 2016)
$\mathbb{E}[R_{T}]=O(K^{2}\log T).$
Implications.
The theorem implies that $\tfrac{\mathbb{E}[R_{T}]}{T}\!\to\!0$ as $T\!\to\!\infty$, i.e., PDO with D-TS asymptotically converges to selecting Copeland-optimal prompts. If a Condorcet winner exists, D-TS converges to that unique prompt; otherwise, it converges to the set of Copeland winners.
@@ 3.2 Efficient Prompt Discovery via Top-Performer Guided Mutation
Coverage-Theoretic Justification.
The efficiency of mutating top performers can be understood by assuming an underlying performance function $C$ (e.g., expected Copeland score) that is $L$-Lipschitz under a prompt distance $d(\cdot,\cdot)$ measuring semantic or functional similarity between prompts. The $\epsilon$-neighborhood of the optimum is $B_{\epsilon}(p^{*})=\{p\in\mathcal{P}:d(p,p^{*})\leq\epsilon\}$.
Drawing on classical results in Lipschitz bandits (Kleinberg et al., 2008), concentrating mutations around the current best prompt is analogous to “zooming in” on near-optimal regions. Mutating a weak prompt $p_{\text{weak}}$ produces only local variants; unless $p_{\text{weak}}$ lies close to $p^{*}$, reaching $B_{\epsilon}(p^{*})$ requires many steps, with expected hitting time increasing in $d(p_{\text{weak}},p^{*})$. By contrast, mutating the best-known prompt $p_{\text{top}}$ yields neighbors that satisfy
$C(p)\;\geq\;C(p_{\text{top}})-L\cdot d(p,p_{\text{top}}),$
so they inherit relatively strong performance. This biases exploration toward high-performing regions and shortens convergence.
PDO thus combines the regret guarantees of D-TS with the coverage-theoretic efficiency of top-performer mutation, making it a sample-efficient approach for prompt optimization without ground-truth references (see Algorithm 1). In practice, we periodically prune the candidate set by removing prompts with low Copeland scores at the current iteration to accelerate convergence.

### v2

@@ Abstract
Large language models (LLMs) are highly sensitive to prompts, but most automatic prompt optimization (APO) methods assume access to ground-truth references (e.g., labeled validation data) that are costly to obtain. We propose the Prompt Duel Optimizer (PDO), a sample-efficient framework for label-free prompt optimization based on pairwise preference feedback from an LLM judge. PDO casts prompt selection as a dueling-bandit problem and combines (i) Double Thompson Sampling to prioritize informative comparisons under a fixed judge budget, with (ii) top-performer guided mutation to expand the candidate pool while pruning weak prompts. Experiments on BIG-bench Hard (BBH) and MS MARCO show that PDO consistently identifies stronger prompts than label-free baselines, while offering favorable quality–cost trade-offs under constrained comparison budgets.
@@
Yuanchen Wu1thanks: Work done during an internship at Meta.thanks: Code is available at https://github.com/meta-llama/prompt-ops   Saurabh Verma2   Justin Lee2   Fangzhou Xiong2 Poppy Zhang2Amel Awadelkarim2Xu Chen2Yubai Yuan1Shawndra Hill2 1Department of Statistics, The Pennsylvania State University 2Meta
@@ 1 Introduction
To address these issues, we introduce the Prompt Duel Optimizer (PDO), which casts label-free prompt optimization as a dueling-bandit problem where the cost is the number of judge comparisons. PDO combines Double Thompson Sampling (D-TS) to focus comparisons on informative prompt pairs with a top-performer mutate-and-prune strategy that expands the pool while discarding weak candidates. We further assess judge reliability via cross-family re-evaluation to probe circularity and show that PDO’s gains are robust across judge choices.
@@
    We formalize label-free prompt optimization with LLM-judged pairwise comparisons as a dueling-bandit problem.
@@
    We propose PDO, integrating D-TS for sample-efficient prompt selection with top-performer guided mutation that steers search toward stronger regions and yields better candidates.
@@
    We show empirically on BBH and MS MARCO that PDO outperforms label-free baselines under fixed judge budgets, and analyze quality–cost trade-offs via budget and runtime comparisons.
@@
    We analyze judge noise and potential circularity, and demonstrate that PDO’s gains persist under alternative judge families.
@@ 3.2 Efficient Prompt Discovery via Top-Performer Guided Mutation
3.3 Theoretical insights of PDO
PDO can be viewed as a two-level optimization scheme that alternates selection and discovery. Within any fixed candidate pool, D-TS provides a principled way to allocate pairwise comparisons and is known to achieve $O(K^{2}\log T)$ expected Copeland regret in general settings, implying vanishing average regret and convergence to the Copeland-winner set as the comparison budget grows (Wu and Liu, 2016).
To move beyond a fixed pool, PDO periodically expands the candidate set by mutating the current Copeland leader. Under a standard smoothness assumption that prompt utility varies gradually under local edits (e.g., $L$-Lipschitz with respect to a prompt distance), this strategy concentrates new candidates in increasingly competitive neighborhoods, analogous to the “zooming-in” intuition in Lipschitz/metric bandits (Kleinberg et al., 2008).
Taken together, PDO separates prompt optimization into (i) sample-efficient identification of the best prompt among current candidates (via D-TS) and (ii) targeted exploration that increases coverage of near-optimal regions (via top-performer mutation), which helps explain its empirical efficiency under limited judge budgets.
@@ 9      Expand $W,N$ with zero row/column for $p_{\text{new}}$;
4 Experiment and Results
4.1 Experimental Setup
Datasets.
In this section, we conduct experiments to evaluate PDO on both closed-ended multiple-choice tasks and open-ended QA tasks. For the multiple-choice setting, we select 16 tasks from Big-Bench-Hard (Suzgun et al., 2022), using accuracy as the evaluation metric. For the open-ended QA setting, we consider four task categories from MS-MARCO (Bajaj et al., 2016), where the final evaluation metric is an integer score between $1$ and $5$, assigned by an LLM judge comparing the model’s output with the ground-truth answer provided by the original dataset.
LLM Judge Design.
For multiple-choice tasks, we require the LLM to produce both an answer and the accompanying reasoning given the instruction prompt. We then apply a dual-judge approach: if two prompts yield different answers, the LLM judge selects the prompt with the correct answer; if they yield the same answer, the judge decides based on the quality of reasoning. For open-ended tasks, the LLM judge design is more straightforward: each pair of responses is evaluated on accuracy, completeness, relevance, and clarity, consistent with the criteria of the final evaluation metric. To mitigate position bias, the order of the two outputs in each prompt pair is randomized before being fed into the judge template. A detailed rationale and analysis of the LLM preference judge design are provided in Section 5 and Appendix E.
(a) Description
(b) Entity
(c) Numeric
(d) Location
Baselines.
For the Big-Bench-Hard (BBH) dataset, we also evaluate PDO under a supervised setting to enable comparison with popular supervised APO methods, including APE (Zhou et al., 2022), OPRO (Yang et al., 2024), and Breeder (Fernando et al., 2023). The optimization itself remains label-free; labels are introduced only in the final stage to select the prompt that achieves the highest development-set accuracy among candidates generated through initial proposals and prompt mutations. Results for the without-label and with-label settings are reported in Tables 1 and 6, respectively.
4.2 Benchmark Results
Multiple-choice Tasks Performance.
As shown in Table 1, using only preference signals from the LLM judge and selecting the winner prompt by Copeland scores, PDO achieves the highest evaluation accuracy on 13/16 tasks. Notable gains include Tracking-7 (0.641 vs. 0.543, +9.8pp) and Web of Lies (0.942 vs. 0.861, +8.1pp).
Open-ended QA Tasks.
5 LLM Judge Analysis
In the experiments reported above, we use two types of LLM judges. The first is the preference judge used during LLM prompt optimization to compare outputs from two candidate prompts. The second is the evaluation judge, which provides the final score for the open-ended MS-MARCO task. In this section, we present additional experiments to analyze the noise and potential bias introduced by both judge types.
5.1 Preference Judge
Correlation Between Judge Noise and Task Performance.
We proxy preference judge effectiveness for the 16 BBH tasks by the test accuracy gap between Table 1 (prompt chosen by Copeland score, label-free) and Table 6 (prompt chosen by development-set accuracy). In our results, Tracking-7 and Web of Lies show small gaps and achieve top performance under both selection criteria, whereas Geometric exhibits a large discrepancy of 11.4 percentage points.
To investigate further, for each task we fix $|\mathcal{P}|=20$ instruction prompts with varying accuracy, run D-TS with the LLM judge, and compare it against an oracle judge that always selects the higher-accuracy prompt in each duel. At round $t$, we record the ground-truth accuracy rank of the current Copeland leader. Figure 3(a) shows that the oracle converges to the best prompt by round 4; with the LLM judge, Tracking-7 steadily improves to rank 2 and Web of Lies approaches rank $\approx 2.5$, while Geometric remains around ranks 6–8 across rounds. These trends confirm that judge reliability is closely related to the performance gaps observed in Tables 1 and 6.
(a) Judge noise across different BBH tasks.
(b) Noise reduction on Geometric using Claude 4.5.
Figure 3: (a) The accuracy of the LLM judge varies across tasks, with some tasks (e.g., Geometric) showing persistent misjudgments compared to an oracle judge. (b) Using a frontier LLM judge (Claude 4.5) helps reduce the impact of noisy judgements.
Reducing Judge Noise using frontier model.
We hypothesize that a key source of judge noise on BBH tasks arises from heterogeneous task difficulty: when two prompts produce different answers, the preference judge must determine which output is correct, and this decision becomes unreliable on harder tasks for judges with limited reasoning ability. A natural remedy is to replace the current Llama-3.3 preference judge with a stronger frontier model that better approximates an oracle, therefore reducing decision noise.
Table 2 reports the selection accuracy of the preference judge on examples where the two candidate answers differ. With Llama-3.3-70B as the judge, accuracy is high on Tracking-7 and Web of Lies but only marginally above chance on Geometric, consistent with the trend in Figure 3(a). Substituting Claude-4.5-Sonnet as the preference judge substantially improves selection accuracy on Geometric and yields near-oracle performance on the other two tasks. As a result, when replacing the preference judge with Claude 4.5, the test accuracy of the prompts on Geometric found by PDO exceeds that of SPO across LLM judge API budgets as reported in Figure 3(b). We report additional results on three other BBH tasks in Figure 7.
Judge Model Geometric Tracking-7 Web of Lies
Llama-3.3-70B 0.59 0.89 0.85
Claude-4.5-Sonnet 0.87 0.98 0.99
Table 2: Preference-judge selection accuracy on BBH instances where two candidate prompts produce different answers (1 = always selects the correct answer; 0.5 = random guessing).
5.2 Evaluation Judge
Cross-family Evaluation.
For the MS MARCO results in Section 4, we use the same Llama 3.3 model both to guide prompt optimization and to score the final prompt, which may introduce judge circularity (i.e., the optimized prompt is favored by the same model used for evaluation). To probe this issue beyond a single prompt pair, we form two prompt sets by sampling uniformly at random: 10 prompts produced by PDO and 10 prompts produced by SPO (all with their original Llama 3.3 judge scores recorded during optimization). To avoid additional inference variability, we use cached model outputs for each prompt on the same evaluation set and re-score these fixed outputs against the reference answers using judges from different model families.
Table 3 reports cross-family re-scoring results across the 10 sampled prompts per method. PDO remains higher than SPO under every judge, with consistently positive margins, suggesting that the observed improvement is not merely an artifact of circularity with the original Llama 3.3 judge.
In Appendix A, we include a small-scale human agreement study testing whether the LLM judge evaluation on MS-MARCO aligns with human preferences on examples where PDO and SPO outputs receive different scores.
Judge Model PDO SPO $\Delta$
Llama 3.3 (original) $4.68\pm 0.06$ $4.50\pm 0.07$ $0.18$
Mistral-Large $61\pm 0.07 4.4$$49\pm 0.07$ $0.12$
GPT-4o $60\pm 0.06 4.4$$41\pm 0.06$ $0.19$
Claude 3.5 Sonnet $4.45\pm 0.08$ $4.25\pm 0.08$ $0.20$
Claude 4.5 Sonnet $58\pm 0.06 4.4$$49\pm 0.06$ $0.09$
Table 3: Cross-family LLM-judge re-evaluation on MS MARCO (Description).
6 Ablation Study
Cost Analysis.
We make two kinds of LLM calls in prompt optimization: inference calls to generate a prompt’s outputs and judge calls to compare outputs pairwise. Because our evaluation is empirical, we assess sample efficiency by comparing methods under the same judge-call budget.
Figure 2 shows that D-TS consistently finds better prompts than random sampling given the same number of judge calls. We also report mean wall-clock time on MS MARCO for SPO with its default 20 rounds and for PDO early-stopped at 10 rounds (Table 4). We use 10 rounds for PDO because Figure 2 indicates that by round 10, the prompt selected by PDO already surpasses SPO in test accuracy.
SPO is faster per run largely because it compares only two prompts per round and immediately eliminates the loser. However, without an explicit exploration–exploitation strategy, this procedure can be brittle under noisy comparisons, which helps explain its weaker performance.
Practically, as discussed in Section 5.1, we control judge noise by using a strong but expensive model (e.g., Claude-4.5) to guide the search, and then deploying the optimized prompt on a lower-cost model (e.g., Llama-3.3) that serves production traffic. In this setting, D-TS is especially effective because it achieves better prompts with fewer expensive judge calls.
Method Description Entity Numeric Location
PDO $\mathbf{4.39}$ $\mathbf{3.39}$ $\mathbf{3.83}$ $\mathbf{3.10}$
SPO 6.78 6.74 6.17 7.45
Table 4: Mean wall-clock time (minutes) on MS MARCO across four prompt tasks.
Prompt Mutation.
PDO incrementally expands the prompt pool by mutating top-performing prompts. We visualize its effect on Web of Lies and Tracking-7. In the mutate variant, at rounds 10 and 20 we prune the 10 lowest–Copeland-score prompts and replace them with 10 new candidates generated by mutating the current Copeland leader. The non_mutate baseline keeps the original 20 prompts fixed throughout. At each round, we report the ground-truth accuracy of the current Copeland leader. As shown in Figure 4, both methods behave similarly before round 10, but mutation yields consistently higher accuracy afterward, surpassing the limits of the fixed pool and navigating toward better prompt regions.
We additionally report results across all 16 BBH tasks in Table 7. In total, 8 out of 16 tasks show statistically significant gains under mutation, and mutation never underperforms relative to the non-mutation baseline.
Pairwise Preference vs. Pointwise Scoring.
7 Related Work
APO Methods: from Supervised to Label-free.
Traditional prompt optimization methods rely heavily on supervised signals from labeled validation sets, where the typical pipeline involves generating, mutating, and scoring prompts Zhou et al. (2022); Fernando et al. (2023); Pryzant et al. (2023); Guo et al. (2023); Schulhoff and others (2024). More recently, studies have reduced supervision requirements Xiang et al. (2025); Chen et al. (2024); Madaan et al. (2023); Cheng et al. (2023); Dong et al. (2022). Notably, Self-Supervised Prompt Optimization (SPO) Xiang et al. (2025) eliminates external references by using output-vs-output comparisons, iteratively generating, executing, and comparing prompts through pairwise LLM judgments. However, SPO follows a greedy hill-climbing loop without a principled exploration–exploitation strategy, limiting its ability to allocate comparisons efficiently and to robustly identify high-performing prompts.
Bandit-Based Prompt Optimization.
8 Conclusion
In this paper, we introduce PDO, a prompt optimization method that reduces dependence on costly supervision by formulating label-free optimization as a dueling bandit problem guided by LLM preference feedback. By combining Double Thompson Sampling (D-TS) with top-performer–guided mutation, PDO adaptively searches for informative prompt comparisons to identify high-performing prompts. Experiments on BBH and MS-MARCO datasets show that PDO consistently achieves competitive performance against baseline methods across tasks. We analyze judge noise and potential circularity, and demonstrate that PDO’s gains persist under alternative judge families.
Limitation
While our dueling-bandit framework provides a practical and sample-efficient approach to prompt optimization without ground-truth labels, it also has inherent limitations. The capability of the preference judge plays a critical role: depending on task complexity, the judge may be noisy. Stronger foundation models tend to yield more reliable preferences but incur substantially higher cost, whereas smaller or less specialized models may introduce greater noise, particularly on domain-specific tasks. Because the algorithm optimizes for the judge’s notion of quality rather than the true task metric, it may favor stylistic patterns that align with the judge’s preferences.
To mitigate these concerns, we include a detailed discussion and ablation studies on LLM judge design, along with a small-scale human evaluation, in Section 5 and Appendix E. However, we were unable to conduct broader experiments across a wider range of tasks. Despite these challenges, our formulation provides a principled and scalable starting point for label-free prompt optimization using preference feedback, and we hope it motivates future work on improving the alignment between LLM judges and task objectives.
Appendix A Small-scale Human Agreement Check
In addition to the cross-judge evaluation in Table 3, we examine whether the original Llama 3.3 MS-MARCO judge aligns with human preferences. We select 20 evaluation examples (5 from each of the four MS-MARCO tasks) for which the cached PDO and SPO outputs receive different Llama 3.3 scores under the 1–5 rubric (e.g., 5 vs. 4). We then recruit 10 non-author volunteer raters (graduate-student participants from a university population) to perform a blinded pairwise comparison: for each example, raters are shown the input and two anonymized outputs (PDO vs. SPO), with left/right order randomized and prompt identities hidden, and are asked to choose the output they prefer under the same criteria as the Llama judge (accuracy, completeness, relevance, and clarity).
For each example, we compute an alignment rate as the fraction of raters whose preferred output matches the direction implied by the Llama scores, and we summarize the distribution of alignment rates across the 20 examples in Figure 6. Overall, this distribution suggests that the Llama-based evaluation signal is broadly consistent with human preferences on the subset of cases most relevant to distinguishing PDO from SPO, with a mean alignment ratio of 0.87.
Figure 6: Histogram of per-example human–Llama alignment rates on 20 MS MARCO examples, where PDO and SPO outputs receive different Llama 3.3 evaluation scores.
Appendix B Analysis of Exploration Strategies
Round Method Description Entity Numeric Location
5 D-TS 4.52 4.38 4.00 4.32
Self-Sparring 4.50 4.40 4.00 4.30
RMED 4.47 4.37 3.94 4.29
RUCB 4.46 4.32 3.92 4.26
MergeRUCB 4.44 4.34 3.88 4.28
Random 4.40 4.36 3.89 4.20
10 D-TS 4.57 4.44 4.07 4.41
Self-Sparring 4.56 4.46 4.06 4.40
RMED 4.53 4.42 4.02 4.39
RUCB 4.52 4.38 4.00 4.36
MergeRUCB 4.50 4.40 3.95 4.38
Random 4.45 4.41 3.97 4.29
15 D-TS 4.61 4.48 4.11 4.45
Self-Sparring 4.59 4.50 4.11 4.43
RMED 4.57 4.45 4.05 4.42
RUCB 4.55 4.40 4.03 4.39
MergeRUCB 4.53 4.42 3.98 4.41
Random 4.48 4.44 4.00 4.33
Table 5: Comparison of exploration strategies on MS-MARCO across optimization rounds in PDO. We report the mean test score of the current Copeland leader over rounds, selected by different exploration strategies. For each {round, dataset}, the best score is shown in bold and the second best is underlined.
The dueling-bandit literature spans several algorithmic families. A useful high-level division is between non-stochastic index policies, which select the next duel deterministically based on confidence or divergence indices, and stochastic policies, which randomize comparisons via posterior sampling. To justify the use of D-TS as the exploration strategy in PDO, we conduct additional experiments comparing two UCB-style index methods (RUCB Zoghi et al. (2013) and MergeRUCB Zoghi et al. (2015)), a KL-divergence index method (RMED Komiyama et al. (2015)), and two posterior-sampling methods (D-TS Wu and Liu (2016) and Self-Sparring Sui et al. (2017)).
Table 5 shows monotonic improvement from Round 5 to 15 for all methods. Under the same judge budget, posterior-sampling approaches are consistently strongest: D-TS achieves the best performance in most (round, task) cells, while Self-Sparring remains highly competitive and is the top-performing method on Entity at every reported round. RMED consistently falls between the TS-based methods and the UCB-style baselines—outperforming RUCB and MergeRUCB but remaining slightly below posterior sampling—while random selection performs worst overall.
We conjecture that this performance gap reflects characteristics of the prompt-optimization setting: a short optimization horizon with noisy preference feedback, where many prompt pairs are difficult to reliably separate and the effective notion of “best” aligns more closely with overall win frequency than with a clearly dominant Condorcet winner. In such regimes, posterior sampling can act as an anytime, uncertainty-aware heuristic that continues to explore multiple plausibly strong prompts without over-investing in resolving near-ties. By contrast, index-based methods (both UCB- and divergence-based) may allocate a larger fraction of a limited budget to shrinking confidence sets or satisfying information constraints, which can be less effective when the horizon is small and the judge signal is noisy.
Appendix C Experiment Details
C.1 Hyperparameters
For the main experiments in Section 4.2, we initialize with 20 candidate prompts for BBH and 50 candidate prompts for MS-MARCO. Each experiment runs for 30 rounds, with 25 duels per round. At rounds 10 and 20, we apply prompt mutation by selecting the top-3 prompts ranked by Copeland scores and generating 10 new prompts. At the same time, we prune the 10 lowest-ranked prompts by Copeland scores. We always select the prompt with the highest Copeland score as the winner, using the average win rate as a tiebreaker when multiple prompts share the same score. The D-TS parameter is fixed at $\alpha=1.2$ throughout all experiments.
C.2 Prompt generation strategies
C.3 Baselines
Unsupervised (label-free) methods
SPO follows an iterative optimize–execute–evaluate loop. In each round, the method first samples $n_{\text{examples}}$ fresh examples and evaluates the current prompt by running one answer-generation LLM call per example. A revised prompt is then proposed using a single LLM call conditioned on the observed cases. Both the current and revised prompts are evaluated on the same set of examples via an additional $n_{\text{examples}}$ answer-generation calls. The two sets of outputs are subsequently compared using $n_{\text{judge\_trials}}$ pairwise LLM judge calls, where each judge observes both outputs with randomized order to mitigate positional bias and votes for the preferred prompt. The revised prompt is accepted if it receives a majority of the judge votes; otherwise, the current prompt is retained. Overall, each round incurs $2n_{\text{examples}}+n_{\text{judge\_trials}}+1$ LLM calls.
With the default setting $n_{\text{examples}}=3$ and $n_{\text{judge\_trials}}=3$, this corresponds to 10 calls per round, and we run 20 rounds to obtain the main results reported in Section 4.
Supervised methods
For supervised APO methods, the typical pipeline involves generating, mutating, and scoring prompts. APE frames prompt design as black-box optimization, generating candidate instructions from input–output demonstrations and refining them through iterative search. OPRO treats the LLM itself as the optimizer: a meta-prompt with task descriptions and prior results guides the model to propose and evaluate new candidates in an iterative loop. Breeder applies an evolutionary approach, jointly evolving task-prompts and mutation-prompts through LLM-driven mutation and selection.
C.4 Datasets
BIG-bench Hard (BBH).
MS MARCO.
Appendix D Discussion on Evaluation Scope
We evaluate PDO on BBH and MS MARCO, which capture two practically important prompt-optimization regimes. BBH covers a broad set of real-world, prompt-based instruction-following classification tasks in a cold-start setting, where labeled data are scarce and costly to obtain at the outset. This setting is increasingly relevant in industry workflows such as content moderation Palla et al. (2025), where many systems are transitioning from fully supervised ML/DL pipelines to prompt-based solutions whose performance is largely determined by the underlying policy prompt.
In contrast, MS MARCO represents a single-turn instruction/QA setting, mirroring scenarios where practitioners optimize a system prompt for a locally deployed LLM application. In such cases, tasks are often open-ended and evaluation criteria can be subjective, making an LLM judge a natural and scalable choice for preference-based comparisons.
We admit that extending PDO to multi-turn interactions, as well as jointly optimizing sets of prompts (e.g., for multi-agent systems), is a valuable and important direction for future work.